OpenStudy (anonymous):

to simplify the root of 84 i understand the steps but im confused as to do you break the 84 as much as you can for example would it be 2 and 42 or28 and 3 21 and 4 so on ...which one do you choose whats the rule I am supposed to follow to figure out the whole answer

OpenStudy (unklerhaukus):

om so break up into as small bits as you can (ie factorize the number into primes)

OpenStudy (anonymous):

my way I look for the greatest perfect square number that is a factor of the given number

OpenStudy (anonymous):

greatest as in this case 2 x 42?

OpenStudy (unklerhaukus):

\[\sqrt(84)\] \[=\sqrt{(2^2\times3\times 7)}\]

OpenStudy (mathteacher1729):

Keep going until you're left with primes. :) So 84 = 2*42 And 2*42 = 2*2*21 and 2*2*21 = 2*2*7*3 And we can go no further. Now sqrt (2*2*7*3) is really sqrt(2^2) times sqrt(7*3) BUT... sqrt*2^2) = sqrt(4) = 2 and sqrt(7*3) = ... well, we can't simplify that. no "number times itself" under the sqrt. So the final simplified answer is just 4 times sqrt(21). hope this helps. :)

OpenStudy (anonymous):